John is now four times as old as Annie was six years ago. How old is Annie today if John is 20 years old?
A)8
B)11
C)12
D)14

Answer by anantha(86)

sol:
present age of Annie=x years
six years ago Annie age=x-6
present age of john=four times as old as Annie was six years ago
present age of john=4*(x-6)
according to the problem present age of john is 20 years
4*(x-6)=20
dividing by 4 on both sides
4*(x-6)/4=20/4
x-6=5
adding 6 on both sides
x-6+6=5+6
x=11
the present age of Annie=11 years